OpenStudy (anonymous):

ABCD is a //gm. & P is the mid point of AB If ar(APCD)=36cm^2 then ar(triangle ABC).

OpenStudy (anonymous):



OpenStudy (anonymous):

is it 24cm^2

OpenStudy (shubhamsrg):

 ar(APCD) = 36 = 1/2 h (x+ 2x) ...(1) ar(ABC) = 1/2 h (2x)